OpenStudy (neycuurose1):

Which answer shows this equation written in standard form? 8 – 1(x – y) = –3x + 5 A. 2x + y = –3 B. 2x – y = –3 C. –4x + y = –3 D. 2x + y = 13

OpenStudy (gnorris):

2x + y = –3

OpenStudy (gnorris):

Hint: distribute 8 - x + y = -3x + 5 +3x +3x 8+2x+y = 5 -8 -8 2x+y =-3

OpenStudy (neycuurose1):

Thanks @gnorris

OpenStudy (gnorris):

you're welcome